I am not sure if I am doing this problem right?
x-24
_______ =13
4

Answer by checkley75(3666)

(x-24)/4=13 now cross multiply
x-24=4*13
x-24=52
x=52+24
x=76 answer.
proof
(76-24)/4=13
52/4=13
13=13